# Single-Pass Document Scanning for Question Answering

### v1

@@ 1 Introduction
The Single-Pass Scanner outperforms state-of-the-art embedding models across 41 long-document QA benchmarks (see Table 1). It also maintains faster speed at processing documents and uses fewer FLOPs than embedding models while achieving higher accuracy (see Table 2). The Single-Pass Scanner generalizes well on very long documents, achieving performance close to GPT-4o’s full-context capabilities on documents longer than 256k tokens (see Figure 3), while only using 1600 tokens from 50 relevant sentences it identifies.
@@ 2 Related Work
State Space Models: Meanwhile, SSMs emerge as an alternative to process long sequences (Fu et al., 2024; Gu et al., 2022; Peng et al., 2023; Arora et al., 2024), as they have linear scaling during training and inference. Dao & Gu (2024) incorporated input-dependent parameters into SSMs and integrate efficient parallelizable training and efficient autoregressive inference. Glorioso et al. (2024) and Waleffe et al. (2024) proposed a hybrid Mamba that combines Mamba with attention. Arora et al. (2024) used non-causal prefix-linear-attention to improve model understanding of the global context. Some recent works built embedding models from SSMs (Hwang et al., 2024; Zhang et al., 2024a), for example, from the Monarch Mixer architecture (Saad-Falcon et al., 2024; Fu et al., 2024).
Retrieval-Augmented Generation (RAG): The approach of retrieving information using an embedding model followed by generating an answer has been fundamental for processing long texts that exceed the context limits of language models (Nakano et al., 2021; Borgeaud et al., 2022; Wang et al., 2023a; Izacard & Grave, 2020; Huang et al., 2023; Liu et al., 2024b; Xu et al., 2024; Yu et al., 2024).
Transformer-based Embedding Models: Transformer-based embedding models are typically used as retrievers for RAG systems. Previous embedding models focus on semantic understanding and instruction-following (OpenAI, 2024; Wang et al., 2023b; Izacard et al., 2021; Lin et al., 2023; BehnamGhader et al., 2024). Embedding models NV-Embed-v2-7B (Lee et al., 2024), GTE-Qwen2 (Li et al., 2023), Stella (Zhang, 2024) and GritLM (Muennighoff et al., 2024) excel at identifying semantically similar sentences within localized contexts (Muennighoff et al., 2023). Since embedding models suffer from the lack of contextual understanding, Morris & Rush (2024) proposed context-aware embedding models.
@@ 3.1 Preliminaries
$\begin{gathered}h_{0}=Bx_{0}\quad\dots\quad\left\{\begin{aligned} h_{t}&=Ah_{t-1}+Bx_{t}\\ y_{t}&=C^{T}h_{t}\end{aligned}\right.\quad\dots\quad\left\{\begin{aligned} h_{T}&=Ah_{T-1}+Bx_{T}\\ y_{T}&=C^{T}h_{T}\end{aligned}\right.\end{gathered}$
@@ 4.1 Link-based Generation
The link-based generation strategy improves on pair-based strategy and produces more realistic and coherent questions by leveraging natural connections between different sections of a document. An LLM is initially presented with one random 20-sentence chunk and is then provided with several hundred tokens of nearby context from the broader document. The model is asked to search for meaningful dependencies, such as the reappearance of a character, event, or theme. Once it identifies a relevant link, it is asked to generate a question that depends on both the initial chunk and the discovered linked section. The LLM then outputs a list of chunks along with their connections to the initial chunk. As illustrated in Figure 2, a synthetic question is generated based on a natural connection, and relevant sentences are labeled. Because it relies on textual relationships rather than purely semantic embedding scores, the link-based method tends to generate more coherent training examples. See Appendix B.3 for prompts used.
GPT-4o-mini-0718 is the LLM used for synthetic data generation. See Table 4 for the financial cost of each strategy. See Appendix B.4 for examples of synthetic data generated from each strategy and human evaluation of these synthetic examples.
@@ 5.2 Evaluated Systems
Full-context LLMs: LLMs such as GPT-4o and Llama 3.1 process the entire document in-context and answer the question directly. We also fine-tuned Mamba-2 for full context answer generation in Appendix E.3.
@@
RAG with embedding models: For fairness, we consider two setups. The “5 chunks” setup follows the standard RAG setup (Xu et al., 2023; Li et al., 2024) where the documents are processed in chunks of fixed-length 300 words, and embedding models retrieve the top 5 chunks. The “50 sentences” setup matches Single-Pass Scanner’s approach. The “5 chunks” setup always achieves higher performance for embedding models and BM25 (See Table 2). Therefore, we only report the “5 chunks” results for both embedding models and BM25 in the main paper for brevity. In Appendix A.4, we also report the “50 sentences” setup for all embedding models for completeness.
@@ 5.3 GPT-4o as Judge
The accuracy of freeform answers is evaluated using GPT-4o-0806, which uses a specialized prompt to compare attempted answers with ground-truth answers, providing a binary “yes” or “no” judgment. The prompt is developed from 100 human-annotated examples. On a separate held-out test set of 180 human-annotated examples, GPT-4o’s 180 yes/no judgments have a high agreement with human judgments, achieving a 0.942 macro F1 score. See Appendix C for the prompt.
@@ 5.4 Sliding Window
Our Single-Pass Scanner can generalize beyond its training context length of 10k tokens (see Section 7.2). For instance, the SPScanner 1.3B can handle up to 256k tokens without memory errors on a single node with 8$\times$ 80GB H100. To ensure a fair comparison with GPT-4o, we use the same sliding window approach for Single-Pass Scanners when documents exceed 120k tokens. This allows both models to operate within the same effective context window. Sentences scored twice have their scores (i.e., logit values) averaged.
@@ 6 Main Results
Single-Pass Scanners outperform embedding models: Table 1 reports model performance grouped by document types and the average performance across all data points, with GPT-4o as generator. Both Single-Pass Scanners outperform BM25, all embedding baselines and fine-tuned embedding models, including MTEB (Muennighoff et al., 2023) leaders (as of January 1, 2025), NV-Embed-v2-7B (Lee et al., 2024) and Stella-1.5B (Zhang, 2024). Results on individual dataset performance are provided in Appendix A.4.
@@
Link-based synthetic data is more suitable to train state-space models: Table 1 shows no improvement on GTE-Qwen-1.5B-FT (i.e., fine-tuned) and Contriever-110M-FT over their pre-trained checkpoints, suggesting Single-Pass Scanners learned more than superficial artifacts such as domain adaptability from training documents.
@@ 7 Ablations
Table 4: Comparison of synthetic data strategies for training Single-Pass Scanner 130M, with GPT-4o as the generator.
@@ 7.1 Comparing Synthetic Data Strategies
A main contribution of this paper is our novel link-based synthetic generation strategy. We now evaluate its effectiveness against the two other baseline strategies, chunk-based and pair-based generations. We train 130M Single-Pass Scanners under identical experimental conditions with 500k synthetic questions created from the same documents by each of the three strategies. Table 4 shows that the link-based strategy achieves the strongest results. Interestingly, pair-based generation strongly underperforms link-based generation, suggesting flaws in its synthetic questions. Refer to Appendix B.4 Table 20 for some flawed synthetic questions generated from the pair-based strategy. By discovering connections between chunks of text in a document, the link-based strategy is able to generate more coherent and contextually relevant questions that would require information from distinct parts of the document.
@@ 7.3 Further Analyses
In Appendix E.2, we investigate the feasibility of using an LLM to generatively select sentences via next token prediction. In Appendix E.3, we evaluate the capability of training Mamba-2 to directly answer questions from full document contexts. Additionally, in Appendix E.4, we found the increasing distance between linked-chunks and the query improves Single-Pass Scanners. In Appendix E.5, we found longer training document length of up to 10k tokens improves Single-Pass Scanners. In Appendix E.6, we observed Single-Pass Scanners performed slightly worse when important information is located at the end of a long document. In Appendix E.1, we analyze model performance on scientific documents.
@@ 8 Conclusion
      Test Sets & Validation Sets Overview A.1
@@
      Test Sets Statistics A.2
@@
      Validation Sets Statistics A.3
@@
    Synthetic Data Generation B
@@
      Chunk-based Generation B.1
@@
      Pair-based Generation B.2
@@
      Link-based Generation B.3
@@
      Synthetic Data Quality Evaluation B.4
@@
      Freeform Question-Answer Judging Prompt C.1
@@
      Multiple Choice Question Question-Answer Judging Prompt C.2
@@
    Training Hyperparameter Setting D
@@
      Model Performance on Scientific Documents E.1
@@
      Discriminative Models vs. Generative Models E.2
@@
      Direct Answer Generation on Full Context E.3
@@
      Ablation for Relative Position of Linked-Chunks E.4
@@
      Ablation for Training Document Length E.5
@@
      Are Single-Pass Scanners Lost in the Middle? E.6
@@
      Retrieval Comparison between Single-Pass Scanner and NV-Embed-v2 E.7
@@ A.1 Test Sets & Validation Sets Overview
Table 5: Document type statistics for the 41 benchmark tasks.
@@
Educational $1967$ 45 1594515945\,15945 159 $1595$ $372$ 1.04 6.39
Creative $1733$ 120 585120585120\,585120 585 $1247$ $486$ 1.00 11.16
Official $1328$ 73 8087380873\,80873 808 $1156$ $172$ 1.11 23.18
Conversational $707$ 36 9223692236\,92236 922 $530$ $177$ 1.13 4.97
Total $5735$ 69 1196911969\,11969 119 $4528$ $1207$ $6061$ 11.67
@@ A.2 Test Sets Statistics
narrativeqa $200$ 30 5513055130\,55130 551
qasper $200$ 5 03950395\,0395 039
multifieldqa_en $147$ 6 95169516\,9516 951
hotpotqa $200$ 12 8021280212\,80212 802
2wikimqa $199$ 7 15271527\,1527 152
musique $200$ 15 5601556015\,56015 560
longbook_choice_eng $200$ 194 984194984194\,984194 984
longdialogue_qa_eng $200$ 109 994109994109\,994109 994
longbook_qa_eng $200$ 195 284195284195\,284195 284
loogle_CR_mixup_16k $99$ 31 6333163331\,63331 633
loogle_CR_mixup_32k $99$ 50 3055030550\,30550 305
loogle_CR_mixup_64k $99$ 96 7509675096\,75096 750
loogle_CR_mixup_128k $99$ 177 463177463177\,463177 463
loogle_CR_mixup_256k $99$ 339 055339055339\,055339 055
loogle_MIR_mixup_16k $139$ 33 2403324033\,24033 240
loogle_MIR_mixup_32k $139$ 49 9914999149\,99149 991
loogle_MIR_mixup_64k $139$ 97 8189781897\,81897 818
loogle_MIR_mixup_128k $139$ 178 771178771178\,771178 771
loogle_MIR_mixup_256k $139$ 340 596340596340\,596340 596
multifieldqa_en_mixup_16k $101$ 28 1942819428\,19428 194
multifieldqa_en_mixup_32k $101$ 52 8105281052\,81052 810
multifieldqa_en_mixup_64k $101$ 101 375101375101\,375101 375
multifieldqa_en_mixup_128k $101$ 197 624197624197\,624197 624
multifieldqa_en_mixup_256k $101$ 390 300390300390\,300390 300
muld_CAC $86$ 48 9934899348\,99348 993
ELITR_Bench_QA $130$ 11 1471114711\,14711 147
altqa_4k $199$ 3 22332233\,2233 223
altqa_16k $199$ 13 0111301113\,01113 011
meetingpred_4k $100$ 3 67636763\,6763 676
meetingpred_16k $100$ 11 6921169211\,69211 692
meetingqa_4k $86$ 2 73127312\,7312 731
meetingqa_16k $19$ 72097209\,7209 720
paperqa_4k $82$ 3 11431143\,1143 114
paperqa_16k $90$ 6 67166716\,6716 671
tpo $200$ 3 55535553\,5553 555
financial_qa $68$ 5 $05$\,0505 050
legal_contract_qa $130$ 25 5292552925\,52925 529
scientific_qa $161$ 4 40544054\,4054 405
quality $200$ 5 $95$\,9595 959
coursera $172$ 8 26982698\,2698 269
docfinQA $200$ 212 751212751212\,751212 751

### v2

@@ 1 Introduction
The Single-Pass Scanner outperforms state-of-the-art embedding models across 41 long-document QA benchmarks (see Table 1). It also maintains faster speed at processing documents and uses fewer FLOPs than embedding models while achieving higher accuracy (see Table 2). The Single-Pass Scanner generalizes well on very long documents, achieving performance close to GPT-4o’s (OpenAI, 2024) full-context capabilities on documents longer than 256k tokens (see Figure 3), while only using 1600 tokens from 50 relevant sentences it identifies.
@@ 2 Related Work
State Space Models: Meanwhile, SSMs emerge as an alternative to process long sequences (Fu et al., 2024; Gu et al., 2022; Peng et al., 2023; Arora et al., 2024), as they have linear scaling during training and inference. Dao & Gu (2024) incorporated input-dependent parameters into SSMs and integrate efficient parallelizable training and efficient autoregressive inference. Glorioso et al. (2024) and Waleffe et al. (2024) proposed a hybrid Mamba that combines Mamba with attention. Arora et al. (2024) used non-causal prefix-linear-attention to improve model understanding of the global context. Some recent works built embedding models from SSMs (Hwang et al., 2024; Zhang et al., 2024a), for example, the M2-BERT (Saad-Falcon et al., 2024) embedding model from the Monarch Mixer architecture (Fu et al., 2024).
Retrieval-Augmented Generation (RAG): The approach of retrieving information using an embedding model followed by generating an answer has been fundamental for processing long texts that exceed the context limits of language models (Nakano et al., 2021; Borgeaud et al., 2022; Wang et al., 2023a; Izacard & Grave, 2020; Huang et al., 2023; Liu et al., 2024b; Xu et al., 2024b; Yu et al., 2024).
Transformer-based Embedding Models: Transformer-based embedding models are typically used as retrievers for RAG systems. Previous embedding models focus on semantic understanding and instruction-following (OpenAI, 2024; Wang et al., 2023b; Izacard et al., 2021; Lin et al., 2023; BehnamGhader et al., 2024). Embedding models NV-Embed-v2-7B (Lee et al., 2024), GTE-Qwen2 (Li et al., 2023), Stella (Zhang, 2024) and GritLM (Muennighoff et al., 2024) excel at identifying semantically similar sentences within localized contexts (Muennighoff et al., 2023). Since embedding models suffer from the lack of contextual understanding, Morris & Rush (2024) proposed context-aware embedding models. Xu et al. (2024a); Huang et al. (2024) utilize transformer-based embeddings in order to shrink the retrieval context provided to downstream LLMs.
@@ 3.1 Preliminaries
[…]
@@ 4.1 Link-based Generation
The link-based generation strategy improves on pair-based strategy and produces more realistic and coherent questions by leveraging natural connections between different sections of a document. An LLM is initially presented with one random 20-sentence chunk and is then provided with a 200-sentence context window that includes the initial chunk at the beginning or the end. The model identifies meaningful dependencies (e.g., the thematic reappearance of a character) of the initial chunk with another chunk inside the window. This process is not recursive. Once it identifies a relevant link and this other chunk, it is asked to generate a question that depends on both chunks. The LLM then outputs a list of chunks along with their connections to the initial chunk. As illustrated in Figure 2, a synthetic question is generated based on a natural connection, and relevant sentences are labeled. Because it relies on textual relationships rather than purely semantic embedding scores, the link-based method tends to generate more coherent training examples. See Appendix B.3 for prompts used.
GPT-4o-mini-2024-07-18 is the LLM used for synthetic data generation. See Table 4 for the financial cost of each strategy. See Appendix B.4 for examples of synthetic data generated from each strategy and human evaluation of these synthetic examples.
@@ 5.2 Evaluated Systems
Full-context LLMs: LLMs such as GPT-4o and Llama 3.1 (AI@Meta, 2024) process the entire document in-context and answer the question directly. We also fine-tuned Mamba-2 for full context answer generation in Appendix E.3.
@@
RAG with embedding models: For fairness, we consider two setups. The “5 chunks” setup follows the standard RAG setup (Xu et al., 2023; Li et al., 2024) where the documents are processed in chunks of fixed-length 300 words, and embedding models retrieve the top 5 chunks. The “50 sentences” setup matches Single-Pass Scanner’s approach. The “5 chunks” setup always achieves higher performance for embedding models (e.g., Dragon (Lin et al., 2023)), and BM25 (Trotman et al., 2014; Robertson & Walker, 1994) (See Table 2). Therefore, we only report the “5 chunks” results for both embedding models and BM25 in the main paper for brevity. In Appendix A.4, we also report the “50 sentences” setup for all embedding models for completeness.
@@ 5.3 GPT-4o as Judge
The accuracy of freeform answers is evaluated using GPT-4o-2024-08-06, which uses a specialized prompt to compare attempted answers with ground-truth answers, providing a binary “yes” or “no” judgment. The prompt is developed from 100 human-annotated examples. On a separate held-out test set of 180 human-annotated examples, GPT-4o’s 180 yes/no judgments have a high agreement with human judgments, achieving a 0.942 macro F1 score. See Appendix C for the prompt.
@@ 5.4 Sliding Window
Our Single-Pass Scanner can generalize beyond its training context length of 10k tokens (see Section 7.2). For instance, the SPScanner 1.3B can handle up to 256k tokens without memory errors on a single node with $\times 8$0GB H100. To ensure a fair comparison with GPT-4o, we use the same sliding window approach for Single-Pass Scanners when documents exceed 120k tokens. This allows both models to operate within the same effective context window. Sentences scored twice have their scores (i.e., logit values) averaged.
@@ 6 Main Results
Single-Pass Scanners outperform embedding models: Table 1 reports model performance grouped by document types and the average performance across all data points, with GPT-4o-2024-08-06 as generator. Both Single-Pass Scanners outperform BM25, all embedding baselines and fine-tuned embedding models, including MTEB (Muennighoff et al., 2023) leaders (as of January 1, 2025), NV-Embed-v2-7B (Lee et al., 2024) and Stella-1.5B (Zhang, 2024). Results on individual dataset performance are provided in Appendix A.4.
@@
Link-based synthetic data is more suitable to train state-space models: Table 1 shows no improvement on GTE-Qwen-1.5B-FT (i.e., fine-tuned) and Contriever-110M-FT over their pre-trained checkpoints, suggesting Single-Pass Scanners learned more than superficial artifacts such as domain adaptability from training documents. The long-range dependencies in link-based data makes it hard for embedding models with short context windows to learn.
@@ 7 Ablations
Table 4: Synthetic data strategies for training Single-Pass Scanner 130M. GPT-4o is generator.
@@ 7.1 Comparing Synthetic Data Strategies
This paper introduced our novel link-based synthetic generation strategy. We now evaluate its effectiveness against the two other baseline strategies, chunk-based and pair-based generations. We train 130M Single-Pass Scanners under identical experimental conditions with 500k synthetic questions created from the same documents by each of the three strategies. Table 4 shows that the link-based strategy achieves the strongest results. Interestingly, pair-based generation strongly underperforms link-based generation, suggesting flaws in its synthetic questions. Refer to Appendix B.4 Table 21 for some flawed synthetic questions generated from the pair-based strategy. By discovering connections between chunks of text in a document, the link-based strategy is able to generate more coherent and contextually relevant questions that would require information from distinct parts of the document.
@@ 7.3 Further Analyses
In Appendix E.2, we investigate the feasibility of using an LLM to generatively select sentences via next token prediction. Results show that generative baselines underperform the discriminative SPScanner. In Appendix E.3, we evaluate the capability of training Mamba-2 to directly answer questions from full document contexts. We found its performance remains substantially lower than that of Single-Pass Scanner. Additionally, in Appendix E.4, we found the increasing distance between linked-chunks and the query improves Single-Pass Scanners. In Appendix E.5, we found longer training document length of up to 10k tokens improves Single-Pass Scanners. In Appendix E.6, we observed Single-Pass Scanners performed slightly worse when important information is located at the end of a long document. In Appendix E.1, we analyze model performance on scientific documents. Model performance on scientific documents are generally consistent with the overall performance.
8 An Information Retrieval Perspective
We can alternatively formulate sentence selection as an information retrieval (IR) problem: given a long document, the goal is to retrieve a small subset of sentences sufficient to answer a query. To construct ground-truth labels for this task, we developed an LLM-assisted annotation–validation pipeline.
In the annotation stage, GPT-4o-mini labeled sentences for relevance. Conditioned on the query and its correct answer, the model processed the document in non-overlapping 20-sentence windows, marking sentences deemed relevant. The full annotation prompt is provided in Appendix C.3.
In the validation stage, GPT-4o attempted to answer the query using only the sentences marked as relevant, without access to the gold answer. If the model reproduced the correct answer, the annotations were retained; otherwise, the document was re-annotated using GPT-4o with an expanded 200-sentence window. These revised annotations were then processed with the same validation procedure. Data points failing both validations were discarded, as manual inspection indicated that they required reasoning over context spans too long for the models to reliably capture.
This process yielded 3,539 validated test instances from the original 5,735. Retrieval performance on this filtered set was evaluated using recall, precision, and nDCG. As shown in Table 5, SPScanner 1.3B consistently achieved the best performance across all metrics.
Model
Metric @k BM25 Dragon GTE-Qwen2 OpenAI NV-Embed SPScanner SPScanner
110M 1.5B v3-Large v2-7B 130M 1.3B
Recall 1 $11.5$ $12.5$ $14.2$ $15.1$ $15.8$ $16.1$ 19.4
5 $27.3$ $28.7$ $31.5$ $35.2$ $36.8$ $39.6$ 45.5
10 $35.2$ $36.8$ $40.6$ $45.3$ $46.3$ $51.3$ 57.1
50 $56.3$ $59.6$ $64.5$ $68.2$ $69.0$ $74.4$ 79.0
Precision 1 $28.3$ $30.7$ $34.0$ $37.4$ $38.7$ $41.0$ 47.8
5 $15.8$ $16.7$ $18.3$ $20.7$ $21.5$ $24.1$ 27.4
10 $11.2$ $11.7$ $12.9$ $14.4$ $14.6$ $17.0$ 19.0
50 $4.4$ $4.7$ $5.0$ $5.3$ $5.4$ $6.2$ 6.6
nDCG 1 $28.3$ $30.7$ $34.0$ $37.4$ $38.7$ $41.0$ 47.8
5 $27.6$ $29.3$ $32.2$ $35.8$ $37.3$ $40.5$ 46.7
10 $30.3$ $31.9$ $35.3$ $39.1$ $40.4$ $44.4$ 50.5
50 $36.7$ $38.8$ $42.5$ $46.1$ $47.3$ $51.6$ 57.3
Table 5: The information retrieval performances of SPScanners and embedding models on a set of 3,539 data points. The information retrieval benchmark is constructed in Section 8.
@@ 9 Conclusion
      Test Sets & Validation Sets Overview .A.1
@@
      Test Sets Statistics .A.2
@@
      Validation Sets Statistics .A.3
@@
    Synthetic Data Generation .B
@@
      Chunk-based Generation .B.1
@@
      Pair-based Generation .B.2
@@
      Link-based Generation .B.3
@@
      Synthetic Data Quality Evaluation .B.4
@@
      Freeform Question-Answer Judging Prompt .C.1
@@
      Multiple Choice Question Question-Answer Judging Prompt .C.2
    3. 3.
      Relevant Sentence Annotation Prompt .C.3
@@
    Training Hyperparameter Setting .D
@@
      Model Performance on Scientific Documents .E.1
@@
      Discriminative Models vs. Generative Models .E.2
@@
      Direct Answer Generation on Full Context .E.3
@@
      Ablation for Relative Position of Linked-Chunks .E.4
@@
      Ablation for Training Document Length .E.5
@@
      Are Single-Pass Scanners Lost in the Middle? .E.6
@@
      Retrieval Comparison between Single-Pass Scanner and NV-Embed-v2 .E.7
    8. 8.
      Performance Change around 128-256k Tokens .E.8
    9. 9.
      SPScanner Qualitative Error Analysis .E.9
@@ A.1 Test Sets & Validation Sets Overview
Educational $1967$ 45 15945\,159 $1595$ $372$ 1.04 6.39
Creative $1733$ 120 585120\,585 $1247$ $486$ 1.00 11.16
Official $1328$ 73 80873\,808 $1156$ $172$ 1.11 23.18
Conversational $707$ 36 92236\,922 $530$ $177$ 1.13 4.97
Total $5735$ 69 11969\,119 $4528$ $1207$ $6061$ 11.67
Table 6: Document type statistics for the 41 benchmark tasks.
@@ A.2 Test Sets Statistics
narrativeqa $200$ 30 55130\,551
qasper $200$ 5 0395\,039
multifieldqa_en $147$ 6 9516\,951
hotpotqa $200$ 12 80212\,802
2wikimqa $199$ 7 1527\,152
musique $200$ 15 56015\,560
longbook_choice_eng $200$ 194 984194\,984
longdialogue_qa_eng $200$ 109 994109\,994
longbook_qa_eng $200$ 195 284195\,284
loogle_CR_mixup_16k $99$ 31 63331\,633
loogle_CR_mixup_32k $99$ 50 30550\,305
loogle_CR_mixup_64k $99$ 96 75096\,750
loogle_CR_mixup_128k $99$ 177 463177\,463
loogle_CR_mixup_256k $99$ 339 055339\,055
loogle_MIR_mixup_16k $139$ 33 24033\,240
loogle_MIR_mixup_32k $139$ 49 99149\,991
loogle_MIR_mixup_64k $139$ 97 81897\,818
loogle_MIR_mixup_128k $139$ 178 771178\,771
loogle_MIR_mixup_256k $139$ 340 596340\,596
multifieldqa_en_mixup_16k $101$ 28 19428\,194
multifieldqa_en_mixup_32k $101$ 52 81052\,810
multifieldqa_en_mixup_64k $101$ 101 375101\,375
multifieldqa_en_mixup_128k $101$ 197 624197\,624
multifieldqa_en_mixup_256k $101$ 390 300390\,300
muld_CAC $86$ 48 99348\,993
ELITR_Bench_QA $130$ 11 14711\,147
altqa_4k $199$ 3 2233\,223
altqa_16k $199$ 13 01113\,011
meetingpred_4k $100$ 3 6763\,676
meetingpred_16k $100$ 11 69211\,692
meetingqa_4k $86$ 2 7312\,731
meetingqa_16k $91$ 9 7209\,720
paperqa_4k $82$ 3 1143\,114
paperqa_16k $90$ 6 6716\,671
tpo $200$ 3 5553\,555
financial_qa $68$ 5 0505\,050
legal_contract_qa $130$ 25 52925\,529
scientific_qa $161$ 4 4054\,405
quality $200$ 5 9595\,959
coursera $172$ 8 2698\,269
docfinQA $200$ 212 751212\,751